# More Sample-Efficient Tuning of Particle Accelerators with Bayesian Optimization and Prior Mean Models

### v1

@@ I Introduction
Bayesian optimization (BO) [2, 3] is a model-based optimization algorithm that has shown particular promise towards solving online accelerator control problems in a variety of contexts [4]. This control algorithm builds Gaussian process models (GPs) [5] that use Bayesian statistics to predict objective function values (and corresponding uncertainties) from scratch using previous measurements of the objective function. As a result, BO algorithms can find solutions in fewer optimization steps, which is especially important for solving problems where evaluating the objective function is expensive.
@@
The convergence speed of BO algorithms can be improved further by leveraging its ability to easily encode prior information about the objective function into the model, instead of building the model from scratch every time the optimization is run. Adding prior information to the GP model used in BO improves the model accuracy, leading to faster convergence to optimal solutions, especially when the parameter space is larger than a few variables. For example, previous work [6, 7] showed that incorporating physics information into the kernel function used to predict cross-correlations between adjacent quadrupole focusing magnet strengths improved convergence time when optimizing free electron laser parameters. In the context of accelerator physics, prior information about the objective function can be gathered from a variety of sources, including basic beam dynamics principles, previous optimization runs, historical data sets, and detailed physics simulations.
In this work, we leverage the ability of GP models to incorporate prior information by combining them with fast-executing neural network (NN) surrogate models of the target systems, trained on historical data and beam dynamics simulations. This expands on previous work in [8, 9, 10]. We first describe how non-constant prior mean functions are incorporated into the GP modeling architecture and demonstrate how this affects predictions with and without training data. We then present a simulated study of how incorporating NN prior models impacts optimization of a simulated version of the Linac Coherent Light Source (LCLS) photoinjector [11]. Finally, we present experimental results from using NN prior models at LCLS and the Argonne Tandem Linear Accelerator System (ATLAS) [12].
@@ III Non-Constant Prior Means
Although standard BO is already a sample-efficient algorithm in many practical applications, more challenging optimization problems like those with a high-dimensional search space can still take prohibitively many steps to reach the desired optimization target. As the dimensionality of the problem increases, so does the amount of data needed to make accurate global predictions of the objective function. However, prior information can easily be encoded into GP models. If there is sufficient prior knowledge about the objective function to build an a-priori model, this initial estimate can serve as a more informative prior mean function for the GP and thereby improve the model’s predictive accuracy and BO convergence speed.
@@
The prior mean function can be represented in a number of different ways, including analytic functions, surrogate models or even detailed physics simulations. As a computationally efficient and flexible approach, here we use neural networks to model the prior mean function. Neural networks can be trained on either experimental or simulation data, and they are generally fully differentiable and have a short inference time, which otherwise can make the numerical optimization of the acquisition function a major limitation. They also allow the incorporation of large amounts of data, e.g. from historical measurements or simulations, and thereby complement the GP model which scales relatively poorly with the size of the data set. By representing the prior mean as a neural network, we thus aim to combine the data efficiency of the GP model with the better computational scaling and ability to represent large amount of previous data provided by NNs.
@@ IV Simulation Studies
In order to better understand the influence of a non-constant prior mean on BO performance, its effects are initially studied in simulations using a surrogate model for the LCLS photoinjector. The beamline is parameterized by scalar parameters that control the laser spot size on the cathode, the laser pulse length, the bunch charge, the phases and amplitudes of two RF accelerating cavities, and the focusing strengths of a number of solenoid, normal quadrupole and skew quadrupole magnets. We use a fully connected NN surrogate model trained on data from the beam dynamics simulation code IMPACT-T [14] to predict five scalar quantities of the beam distribution (beam sizes and transverse emittances) at an optical transition radiation diagnostic screen downstream of the photoinjector (OTR2). This surrogate model is incorporated into the GP model built by Xopt [15] as a prior mean function using the LUME-model [16] framework to perform calibration and scaling to experimental units. A detailed description of the surrogate model can be found in Appendix B.
@@ V LCLS Demonstration
The LCLS surrogate model trained on data generated by the LCLS IMPACT-T model was then used as an NN prior mean to perform online optimization of the LCLS photoinjector beamline. We repeatedly performed BO with the EI acquisition function starting with a fixed set of three randomly chosen initial points, with and without the NN prior mean function, to minimize the objective function in Eq. 9. During this, the laser spot size, pulse length, RF cavity parameters and bunch charge were fixed, resulting in a 9-dimensional optimization problem. Results from these optimization runs can be seen in Fig. 4. We observe that using the NN prior model consistently outperforms the constant prior in the first optimization steps. It takes the constant prior five steps longer on average than the NN prior to reach the same optimum. However, after approximately ten steps, the constant prior model surpasses the optimum achieved by the NN prior model, likely due to its low correlation with the measured objective (, ). The NN prior mean model provides an advantage in speed during the early stages of coarse optimization, and it is then surpassed in later refinement of the optimum by standard BO. Based on the simulation results in section IV, it is expected that NN models that are better calibrated to experimental data will increase the performance of BO in this context. In the meantime, it is possible to improve the optimization performance at later steps by transitioning from the NN model to a constant prior mean during the optimization, as discussed in Appendix A.
@@ VI ATLAS Demonstration
We began with establishing the baseline performance of standard BO with a constant prior mean, initiating the optimization with only one data point, resulting in an initial transmission close to , and employed the Upper Confidence Bound (UCB) acquisition function with . This algorithm was executed on the machine using an O beam times, each time starting from the same conditions. Refer to Fig. 5, where the blue dashed line represents the median maximum transmission at each step of the constant prior BO optimization, with the surrounding area depicting the confidence level.
We then trained a neural network model on a dataset comprising experimental data points obtained from a prior experiment involving a N heavy-ion beam to predict transmission as a function of the quadrupole configurations, details of which are discussed in Appendix B. Doing this, we demonstrate and take advantage of effective transfer learning, where historical data gathered during ATLAS operations using one species, N, is used to inform optimization of accelerator operations for a different but similar heavy-ion species, O.
The trained NN model was then incorporated into the BO algorithm as the prior mean and used for optimization with the O beam, employing the same initial conditions as the constant prior BO optimization mentioned earlier. This process was repeated times as well. The results, depicted in orange in Fig. 5, clearly illustrate the superior performance of BO with a neural network prior mean. This approach outperforms traditional BO methods with no prior knowledge, underscoring not only the advantage of incorporating a prior mean model in Gaussian processes but also the potential of transfer learning techniques in optimizing particle accelerators by using models trained under different beams and conditions.
@@ VII Conclusion
Incorporating NN based priors into GP modeling is also a path toward enabling facility-scale BO with hundreds to thousands of input parameters. Neural network surrogate models are able to ingest extremely large sets of historical and simulated data without suffering from an increase in prediction cost, unlike GP models, which scale poorly () with the number of training points. By using NN based surrogate models as prior means we can incorporate much more information into BO about the objective function than we would be able to otherwise. Additionally, combining NN surrogates with GP modeling improves the applicability of using NNs in situations where the NN model does not perfectly predict the real objective function. Gaussian process models can quickly adapt predictions to new data, whereas NN surrogates need significant amounts of data and training time to be calibrated to experimental conditions. Combing the two model types leverages the advantages of both modeling strategies and can be used to provide accurate, facility-scale models of objective functions that are updated in real-time.
Future work in this area should focus on studying metrics for training prior models that yield good performance in the context of BO, including disentangling the effects of the two metrics identified in this work to describe model quality, i.e., MAE and correlation. Additionally, as discussed in Appendix A, it is possible to recover optimization performance on-the-fly, i.e., during optimization, when the prior mean model is a poor predictor of the objective function; future work should more thoroughly explore different heuristics for weighting the use of a NN prior mean function and trainable constant prior functions to improve performance. Finally, non-constant prior mean functions can be incorporated into GP models with constraining functions for use in constrained optimization (see [4] for details) to reduce the number of constraint violations during optimization.
@@ Acknowledgements.
with and as the number of steps increases. Intuitively, one might think about this as a continuous “flattening” of the prior mean function as more steps are taken until it eventually becomes fully constant. Based on the general shape of the performance curves in Fig. 3, we linearly decrease the weight from an initial to at step . As shown in Fig. 7 (green line), this can improve the robustness of Bayesian optimization with a non-constant prior mean, bringing performance closer to that with a constant prior. To take this idea one step further, we let the weighting factor be chosen based on the correlation coefficient determined across previous samples
@@ Appendix B Neural Network Models
The primary objective of implementing this model was to demonstrate that a simple approach for the prior mean could yield benefits in Bayesian optimization, as detailed in this paper. The model was trained using a dataset comprising experimental data points obtained from a prior experiment involving a N ion beam.
Subsequently, the trained model was deployed and evaluated on a different ion species, O, for another experiment. This process showcased the model’s capability to effectively transfer knowledge between different ion beams, highlighting its versatility and generalizability. Different models were trained to different levels of fidelity by conducting the training up to various epochs. The respective model fidelity is reflected in the correlation coefficient and MAE listed in Table 2.
Appendix C Implementations based on PyTorch
Xopt is a Python package that relies on PyTorch to implement the required GP models. As those models are instances of a PyTorch module (torch.nn.Module) they register all trainable sub-modules by default. One benefit of this is that custom parameters can easily be optimized as part of the GP model fit by simply activating their gradients. Yet, as the full GP model is set in training mode before the fit, this will also affect all sub-modules, including the NN prior model. If the NN architecture includes layers like dropout or other modules which change behavior based on whether the model is in training or evaluation mode, this can also affect the prior mean predictions. In the work presented here, this initially led to noisy prior mean predictions and consequently poor BO performance. After its identification, the issue was addressed in Xopt where, by default, prior mean inference is now always performed in evaluation mode.

### v2

@@ I Introduction
Bayesian optimization (BO) [2, 3] is a model-based optimization algorithm that has shown particular promise towards solving online accelerator control problems in a variety of contexts [4]. This control algorithm builds Gaussian process models (GPs) [5] that use Bayesian statistics to predict objective function values (and corresponding uncertainties) from scratch using previous measurements of the objective function. As a result, BO algorithms can find solutions in a sample-efficient manner, which is especially important for solving problems where evaluating the objective function is expensive.
@@
The convergence speed of BO algorithms can be improved further by leveraging its ability to easily encode prior information about the objective function into the model, instead of building the model from scratch every time the optimization is run. Adding prior information to the GP model used in BO improves the model accuracy, leading to faster convergence to optimal solutions, especially when the parameter space is larger than a few variables. In the context of accelerator physics, this type of information can be gathered from a variety of sources, including basic beam dynamics principles, previous optimization runs, historical data sets, and detailed physics simulations.
A straightforward method for incorporating prior information into the GP model is to add available measurement or simulation data directly to the model. This can also be done using multi-task GP models [6], which utilize learned correlations between prior data and measured values to improve the accuracy of model predictions. However, both of these methods suffer from the poor computational scaling of GP models (), preventing the use of large training data sets for online modeling applications.
An alternative method for incorporating physics information into the GP model is to specify a physics-informed kernel function, as demonstrated in [7, 8]. Incorporating physics information into the kernel function allowed the GP model to reduce the data needed to accurately predict cross-correlations between adjacent quadrupole focusing magnet strengths, resulting in improved convergence time when optimizing free electron laser parameters. However, this method is limited to defining correlations between function values at different points in input space, not the absolute value of the function.
In this work, we demonstrate and evaluate a method for incorporating prior physics information into GP models by combining them with fast-executing neural network (NN) surrogate models of the target systems, trained on historical data and beam dynamics simulations. We expand on previous work in [9, 10, 11] to show that combining approximate or well-calibrated NN models with GP models is an effective, computationally inexpensive way to incorporate detailed physics information into BO algorithms, significantly reducing the number of optimization steps needed to converge to a solution. We discuss suitable metrics for model accuracy and the systematic dependency of BO performance on model quality. We also developed software tools, which we describe briefly in Appendix C, to facilitate easy incorporation of prior models into the optimization package Xopt [12], making this technique more broadly accessible to the accelerator community.
We present a simulated study of how incorporating NN prior models impacts optimization of a simulated version of the Linac Coherent Light Source (LCLS) photoinjector [13]. Finally, we present experimental results from using NN prior models trained on simulation at the Linac Coherent Light Source (LCLS) and use NN models to perform transfer learning from one beam species to another at the Argonne Tandem Linear Accelerator System (ATLAS) [14].
@@ III Non-Constant Prior Means
Although standard BO is generally a sample-efficient algorithm in many practical applications, more challenging optimization problems like those with a high-dimensional search space can still take prohibitively many steps to reach the desired optimization target. As the dimensionality of the problem increases, so does the amount of data needed to make accurate global predictions of the objective function. However, prior information can easily be encoded into GP models. If there is sufficient prior knowledge about the objective function to build an a-priori model, this initial estimate can serve as a more informative prior mean function for the GP and thereby improve the model’s predictive accuracy and BO convergence speed.
@@
The prior mean functions can be represented in a number of different ways, including analytic functions, surrogate models or even detailed physics simulations. Ideal models for prior mean function specification are fast executing and are differentiable, i.e., they efficiently compute derivatives of function values with respect to input parameters [16]. These properties are important for fast optimization of the BO acquisition function to select future input parameters.
One such surrogate model which has these properties is a neural network (NN). Neural networks can be trained using a variety of information sources including historical experimental data, simulation data, and physics principles. Additionally, NNs are inherently differentiable owing to the need for gradient calculation for effective model training.
Finally, a critical aspect of NN models is their relatively short evaluation time, which is independent of the training data set size. This enables incorporating extremely large amounts of data into the model from historical measurements or simulations, thereby circumventing the poor computational scaling of GP models. This is demonstrated in Fig. 2, which compares the computational cost of evaluating GP models of the LCLS photoinjector (see Sec. IV) on a fixed number of points with different methods of incorporating prior data into the GP model. If we were to directly incorporate large training data sets into the GP model, the evaluation time scales as . However, if we use a neural network to represent the prior information in the data set the evaluation time of the GP model is substantially smaller and independent of the number of training points. This computational scaling makes a significant practical difference when using these algorithms for online optimization, especially at user facilities that are in high-demand and need to spend as little time as possible tuning the beam.
In practice, and in the cases described here, the NN surrogate model is trained prior to performing optimization using BO. While this does require upfront effort to construct and train a NN surrogate model, the reduction in evaluation time for optimizing high-dimensional optimization problem makes up for the additional work, especially if the optimization is performed repeatedly. Furthermore, NN surrogate models of accelerators are already generated for various applications such as virtual diagnostics [17], virtual accelerator models for online operator feedback [18], and accelerator design [19]. By representing the prior mean as a neural network, we combine the data efficiency of the GP model with the better computational scaling and ability to represent large amount of previous data provided by NNs.
@@ IV Simulation Studies
In order to better understand the influence of a non-constant prior mean on BO performance, its effects are initially studied in simulations using a surrogate model for the LCLS photoinjector. The beamline is parameterized by scalar parameters that control the laser spot size on the cathode, the laser pulse length, the bunch charge, the phases and amplitudes of two RF accelerating cavities, and the focusing strengths of a number of solenoid, normal quadrupole and skew quadrupole magnets. We use a fully connected NN surrogate model trained on data from the beam dynamics simulation code IMPACT-T [20] to predict five scalar quantities of the beam distribution (beam sizes and transverse emittances) at an optical transition radiation diagnostic screen downstream of the photoinjector (OTR2). This surrogate model is incorporated into the GP model built by Xopt [12] as a prior mean function using the LUME-Model [18] framework to perform calibration and scaling to experimental units. A detailed description of the surrogate model can be found in Appendix B, practical details of incorporating the prior mean model into the GP are discussed in Appendix C.
@@ V LCLS Demonstration
The LCLS surrogate model trained on data generated by the LCLS IMPACT-T model was then used as an NN prior mean to perform online optimization of the LCLS photoinjector beamline. We repeatedly performed BO with the EI acquisition function starting with a fixed set of three randomly chosen initial points, with and without the NN prior mean function, to minimize the objective function in Eq. 9. During this, the laser spot size, pulse length, RF cavity parameters and bunch charge were fixed, resulting in a 9-dimensional optimization problem. Results from these optimization runs can be seen in Fig. 5. We observe that using the NN prior model consistently outperforms the constant prior in the first optimization steps. It takes the constant prior five steps longer on average than the NN prior to reach the same optimum. However, after approximately ten steps, the constant prior model surpasses the optimum achieved by the NN prior model, likely due to its low correlation with the measured objective (, ). The NN prior mean model provides an advantage in speed during the early stages of coarse optimization, and it is then surpassed in later refinement of the optimum by standard BO. Based on the simulation results in Sec. IV, it is expected that NN models that are better calibrated to experimental data will increase the performance of BO in this context. In the meantime, it is possible to improve the optimization performance at later steps by transitioning from the NN model to a constant prior mean during the optimization, as discussed in Appendix A.
@@ VI ATLAS Demonstration
We began with establishing the baseline performance of standard BO with a constant prior mean, initiating the optimization with only one data point, resulting in an initial transmission close to , and employed the Upper Confidence Bound (UCB) acquisition function with . This algorithm was executed on the machine using an 16O beam times, each time starting from the same conditions. Refer to Fig. 6, where the blue dashed line represents the median maximum transmission at each step of the constant prior BO optimization, with the surrounding area depicting the confidence level.
We then trained a neural network model on a dataset comprising experimental data points obtained from a prior experiment involving a 14N heavy-ion beam to predict transmission as a function of the quadrupole configurations, details of which are discussed in Appendix B. Doing this, we demonstrate and take advantage of effective transfer learning, where historical data gathered during ATLAS operations using one species, 14N, is used to inform optimization of accelerator operations for a different but similar heavy-ion species, 16O.
The trained NN model was then incorporated into the BO algorithm as the prior mean and used for optimization with the 16O beam, employing the same initial conditions as the constant prior BO optimization mentioned earlier. This process was repeated times as well. The results, depicted in orange in Fig. 6, clearly illustrate the superior performance of BO with a neural network prior mean. This approach outperforms traditional BO methods with no prior knowledge, underscoring not only the advantage of incorporating a prior mean model in Gaussian processes but also the potential of transfer learning techniques in optimizing particle accelerators by using models trained under different beams and conditions.
@@ VII Conclusion
Incorporating NN based priors into GP modeling is also a path toward enabling facility-scale BO with hundreds to thousands of input parameters. Neural network surrogate models are able to ingest extremely large sets of historical and simulated data without suffering from an increase in prediction cost, unlike GP models, which scale poorly with the number of training points. By using NN based surrogate models as prior means we can incorporate much more information about the objective function into BO than we would be able to otherwise. Additionally, combining NN surrogates with GP modeling improves the applicability of using NNs in situations where the NN model does not perfectly predict the real objective function. Gaussian process models can quickly adapt predictions to new data, whereas NN surrogates need significant amounts of data and training time to be calibrated to experimental conditions. Combing the two model types leverages the advantages of both modeling strategies and can be used to provide accurate, facility-scale models of objective functions that are updated in real-time.
Future work in this area should focus on studying metrics for training prior models that yield good performance in the context of BO, including disentangling the effects of the two metrics identified in this work to describe model quality, i.e., MAE and correlation. Additionally, as discussed in Appendix A, it is possible to recover optimization performance on-the-fly, i.e., during optimization, when the prior mean model is a poor predictor of the objective function; future work should more thoroughly explore different heuristics for weighting the use of a NN prior mean function and trainable constant prior functions to improve performance. Furthermore, differentiable beam dynamics models, such as those implemented in Bmad-X [22] or Cheetah [23, 24] should also start to be incorporated into GP models as physics-informed priors. Finally, non-constant prior mean functions can be incorporated into GP models with constraining functions for use in constrained optimization (see [4] for details) to reduce the number of constraint violations during optimization.
@@ Acknowledgements.
with and as the number of steps increases. Intuitively, one might think about this as a continuous “flattening” of the prior mean function as more steps are taken until it eventually becomes fully constant.
Based on the general shape of the performance curves in Fig. 4, we linearly decrease the weight from an initial to at step . As shown in Fig. 8 (green line), this can improve the robustness of Bayesian optimization with a non-constant prior mean, bringing performance closer to that with a constant prior. To take this idea one step further, we let the weighting factor be chosen based on the correlation coefficient determined across previous samples
@@ Appendix B Neural Network Models
The primary objective of implementing this model was to demonstrate that a simple approach for the prior mean could yield benefits in Bayesian optimization, as detailed in this paper. The model was trained using a dataset comprising experimental data points obtained from a prior experiment involving a 14N ion beam.
Subsequently, the trained model was deployed and evaluated on a different ion species, 16O, for another experiment. This process showcased the model’s capability to effectively transfer knowledge between different ion beams, highlighting its versatility and generalizability. Different models were trained to different levels of fidelity by conducting the training up to various epochs. The respective model fidelity is reflected in the correlation coefficient and MAE listed in Table 2.
Appendix C Practical Implementation of Neural Network Priors for Online Control
While the mathematics of incorporating neural networks into GP models is straightforward, the practical implementation in the context of online accelerator control can be challenging. For example, as discussed in Appendix B, the LCLS NN surrogate model was trained using simulation data and there is a significant mismatch between the names and units of the simulation parameters and those of the control parameters at the real machine. Translating the inputs and outputs of this model required a number of transformation layers, shown in Fig. 9, to map control parameters adjusted during optimization to NN input parameters. To address this, we implemented a wrapper class in the LUME-Model package [18] which stores the input and output ordering and the required transformations to make the model applicable to the real machine.
Additionally, specialized modules are needed to enable usage of these surrogate models in the Bayesian optimization package BoTorch [26]. To improve modeling robustness, BoTorch applies data normalization and standardization transformations to input and output data respectively. As a result, incorporating surrogate-based prior mean functions into the GP models created by BoTorch requires specialized modules to un-transform the input data before passing it to the surrogate model and transforming the output data. These modules are built into the Xopt [12] package for convenient use.
Finally, as the models used in this work are instances of a PyTorch module (torch.nn.Module) they register all trainable sub-modules by default. One benefit of this is that custom parameters can easily be optimized as part of the GP model fit by simply activating their gradients. However, as the full GP model is set in training mode before the fit, this will also affect all sub-modules, including the NN prior model. If the NN architecture includes layers like dropout or other modules which change behavior based on whether the model is in training or evaluation mode, this can also affect the prior mean predictions. To address this potential pitfall, modules used as prior mean functions in Xopt are, by default, always queried in evaluation mode.
By creating these open-source software solutions to facilitate easy incorporation of prior models into Bayesian optimization in a reliable way, we are helping to make this technique more broadly accessible to the accelerator community. This reduces development overhead for new facilities and makes the algorithm more ready for use out-of-the-box by accelerator facilities, especially for the growing number that already have available neural network models of their systems.